OpenStudy (anonymous):

the sales of a certain product after an inital release can be found by the equation s=17sqrt4t+61

OpenStudy (mathmale):

Yes? What is the goal of this problem? What's your aim?

OpenStudy (anonymous):

s represents the total sales and t represents the time in weeks after release. How many weeks will pas before the product sells about 125 units?

OpenStudy (mathmale):

\[s=17\sqrt4t+61.or.s=17\sqrt(4t+61)?\] Which one did you mean?

OpenStudy (unklerhaukus):

sub in s= 125, and solve for t

OpenStudy (anonymous):

the first one s = 17sqrt 4t) +61

OpenStudy (unklerhaukus):

? \[s=12\sqrt{4t}+61\]

OpenStudy (anonymous):

instead of 12 it's 17

OpenStudy (unklerhaukus):

\[s=17\sqrt{4t}+61\] \[s=125\] \[125=17\sqrt{4t}+61\]

OpenStudy (anonymous):

so then would I subtract 61 from both sides?

OpenStudy (unklerhaukus):

yes!,

OpenStudy (mathmale):

Your goal is to determine the time, t.

OpenStudy (anonymous):

once I have that it's \[64=17\sqrt{4t}\]

OpenStudy (unklerhaukus):

good work, now divide both sides by 17 ,

OpenStudy (anonymous):

I got \[3.76=\sqrt{4t}\]

OpenStudy (unklerhaukus):

what's the next step to get rid of that √ ?

OpenStudy (anonymous):

square both sides?

OpenStudy (unklerhaukus):

Yes, square both sides !

OpenStudy (anonymous):

divide by 4 and then that would be my answer?

OpenStudy (unklerhaukus):

yeah thats right , you should be left with ... = t [in weeks]

OpenStudy (anonymous):

I got 3.54, but I rounded that up so it's 4 weeks

OpenStudy (unklerhaukus):

The question specifically asks 'How many weeks will pass before 125' units are sold, so i think you have to round down instead